OpenStudy (anonymous):

How do I evaluate the following indefinite integral using u-substitution? cos(5x)/e^sin(5x) dx

OpenStudy (lυἶცἶ0210):

Have you tried sin(5x) as a sub?

OpenStudy (anonymous):

so far I have picked sin(5x) to be my u and then I get cos(5x)5 dx for my du and then I multiply 1/5 to get rid of the 5

OpenStudy (lυἶცἶ0210):

Actually, maybe changing it to something like this might help: \(\LARGE \int ~e^{-sin(5x)} * cos(5x) dx\)

OpenStudy (lυἶცἶ0210):

Now we can sub in -sin5x instead: \( u=-sin(5x) \) and \(du=-5cos(5x) \)

OpenStudy (anonymous):

oh ok I see and I actually tried that but my negative was in the wrong place that might be why I'm struggling

OpenStudy (lυἶცἶ0210):

Yea, after the u sub we end up with this little thing :P \(\LARGE -\frac{1}{5} \int e^u du \)

OpenStudy (lυἶცἶ0210):

Think you can finish it from here?

OpenStudy (anonymous):

yes thank you very much :D

OpenStudy (lυἶცἶ0210):

You're welcome, and welcome to Openstudy btw :) \(\Huge \color{red}{\star^{\star^{\color{green}{\star}}}} \)

OpenStudy (anonymous):

Thanks (: